OpenStudy (anonymous):

Find the height of a square pyramid with volume 126.97 ft³ and dimensions of base are 6 ft by 6 ft

OpenStudy (anonymous):

volume = 1/3 * side^2 * height 126.97 = 1/3 * 36 * height ~